OpenStudy (ttop0816):

Please help! The following system of equations is _____. y = -1/3x + 2/3 2x + 6y = 4 is it inconsistent?? independent equivalent inconsistent

OpenStudy (tkhunny):

Have you considered putting both equations into standard form, like the second one, and comparing them?

OpenStudy (ttop0816):

@tkhunny yes i got 1/3x + y= 2/3

OpenStudy (tkhunny):

That's a good start. Multiply it by 3. "Standard Form" usually asks for integer coefficients if it is possible.

OpenStudy (ttop0816):

then it would be x+3y=2?

OpenStudy (anonymous):

Yes. Remember that any equations that are multiples of each other are equal.

OpenStudy (tkhunny):

Okay, x + 3y = 2 2x + 6y = 4 Why not make your life simpler and remove the common factor from the second equation? Divide everything by 2.

OpenStudy (ttop0816):

oh then it will be parallel!! inconsistent right??

OpenStudy (tkhunny):

No. These two are parallel and inconsistent. x + 3y = 2 x + 3y = 3 There are no solutions. These two are EXACTLY the Same. This makes them consistent, but annoying. x + 3y = 2 x + 3y = 2 There are infinitely many solutions.

OpenStudy (ttop0816):

ohhhh darn ): thanks for the help though! do you mind answering another question for me?

OpenStudy (tkhunny):

gtg to my real job. Sorry. @slickwilly might still be around.